40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1106     Entry time: Sun Nov 2 21:37:22 2008
Author: rana 
Type: Update 
Category: PEM 
Subject: Ranger recovery 
The ranger signal has been bad since around 11 AM on Oct 25 (last Saturday). There are no elog
entries from that day, but I am quite sure that someone must have been working around the PSL
rack area.

It looks like what happened is that someone moved the chair with the monitor on it and/or the wooden
stool next to it. That put tension on the cable connecting the SR560 and the seismometer. The SR560
connector now seems loose and I think probably the cable ground wasn't connected. I swapped the
cable over to the "B" side of the SR560 and the ranger signal is now reasonable (very small offset
and normal seismic signal).

Please be careful when working around there. Everyone always says "I didn't do anything" or "it doesn't
effect anything".

We need to clean up the cabling around there in addition to running a new power cable for the RF amplifier
on the POY table.


I have also reduced its sample rate from 2048 to 512 Hz. The data are OK after 909640694.

I also increased the sample rate of AS_MIC from 2048 to 16384 Hz but that one seems to be broken
---->> the microphone seems to be either disconnected or broken.
ELOG V3.1.3-